OpenStudy (anonymous):

Medal and follow if you can help me solve this. (cos θ + cos θ)^2 + (cos θ + cos θ)^2 a)sin2 θ b)4 c)8cos2 θ d)cos2 θ

OpenStudy (jhannybean):

\[(\cos(\theta) +\cos(\theta))^2 +(\cos(\theta)+\cos(\theta))^2= (2\cos(\theta))^2+(2\cos(\theta))^2\]

OpenStudy (anonymous):

Ah okay so in my head i just got 8cos2 θ

OpenStudy (anonymous):

is that correct in my math? I normally never make calculation mistakes

OpenStudy (jhannybean):

yeah that's the correct answer, but can you explain to me how you got it?

OpenStudy (anonymous):

I just did it in my head. Distributed it. Maybe i'm lucky?